# SSUMamba: Spatial-Spectral Selective State Space Model for Hyperspectral Image Denoising

### v1

@@ Abstract
Denoising hyperspectral images (HSIs) is a crucial preprocessing procedure due to the noise originating from intra-imaging mechanisms and environmental factors. Utilizing domain-specific knowledge of HSIs, such as spectral correlation, spatial self-similarity, and spatial-spectral correlation, is essential for deep learning-based denoising. Existing methods are often constrained by running time, space complexity, and computational complexity, employing strategies that explore these priors separately. While the strategies can avoid some redundant information, considering that hyperspectral images are 3-D images with strong spatial continuity and spectral correlation, this kind of strategy inevitably overlooks subtle long-range spatial-spectral information that positively impacts image restoration. This paper proposes a Spatial-Spectral Selective State Space Model-based U-shaped network, termed Spatial-Spectral U-Mamba (SSUMamba), for hyperspectral image denoising. We can obtain complete global spatial-spectral correlation within a module thanks to the linear space complexity in State Space Model (SSM) computations. We introduce an Alternating Scan (SSAS) strategy for HSI data, which helps model the information flow in multiple directions in 3-D HSIs. Experimental results demonstrate that our method outperforms several compared methods. The source code will be available at https://github.com/lronkitty/SSUMamba.
@@
Hyperspectral image denoising, mamba, global spatial-spectral correlation, deep learningI Introduction
@@
Deep learning-based denoising methods have recently garnered undivided attention. By leveraging data-driven learning, these methods can harness the 3-D nature of HSIs with spectral and spectral information. Due to the challenges posed by the design of methodology, running time, and spatial and computational complexity, existing deep learning-based methods often decompose this 3-D nature into several distinct domains for exploration, such as spatial self-similarity, spectral correlation, and spatial-spectral correlation. For example, recurrent neural networks (RNNs) resort to convolutional neural networks (CNNs) to exploit spectral correlation and spatial-spectral correlation separately [11]. At the same time, attention mechanisms are concatenated to CNNs to capture spatial self-similarity further [12]. Transformer-based methods are also employed to capture long-range correlation in HSIs [9, 13, 13]. However, suffering from the complexity of the self-attention mechanism, these methods incorporate multiple modules to explore different domain knowledge to reduce the length of tokens and avoid redundant information.
@@
However, given the tri-dimensional nature of HSIs, which demonstrate spatial solid continuity and spectral correlation, it is essential to explore the complete global spatial-spectral correlation, especially in image restoration tasks where detailed information tends to degrade.. Subtle long-range spatial-spectral information has a positive impact that should be considered. The State Space Model (SSM) [14, 15], an innovative approach rooted in control systems, has attracted attention for its linear complexity in handling long input sequences. Recently, Mamba [14], an optimization for SSM, has shown unique advantages in computer vision tasks [16, 17], while its linear complexity should offer promising advantages in global spatial-spectral correlation modeling in HSIs. However, how to apply the SSM to HSIs denoising to capture the global spatial-spectral correlation following the 3-D nature of HSIs remains an open question.
@@
This paper introduces a Spatial-Spectral Selective SSM-based HSI denoising method, Spatial-Spectral U-Mamba (SSUMamba), for HSI denoising. To address the disparity between image data and sequence data, we introduce the Vision Mamba (VMamba), aimed at improving the exploration of local texture in the data and mitigating the unidirectional dependency issue present in the SSM model. Our architecture incorporates a Spatial-Spectral Alternating Scan (SSAS) strategy for HSI data, which helps model the information flow in multiple directions in HSIs. Furthermore, we apply Vision Mamba (VMamba) blocks with SSAS strategy in a U-shape network architecture to exploit the multi-scale features to recover the HSI. Thanks to the linear complexity, the SSUMamba can explore the complete global spatial-spectral correlation, and achieves better performance against the compared methods. The contribution of this article can be summarized as follows.
@@
To tackle the difference between image and sequence data, we introduce Vision Mamba (VMamba) block to the HSI modeling to enhance local texture exploration and avoid unidirectional dependency in the SSM model.
The Spatial-Spectral Alternating Scan (SSAS) strategy allows VMamba blocks to exploit global spatial-spectral correlation in all directions, enhancing the 3-D characteristics of HSIs.
Building upon the VMamba blocks with SSAS, we propose a multi-scale SSM-based network, Spatial-Spectral U-Mamba (SSUMamba), for HSI denoising. Experiments demonstrate that SSUMamba achieves state-of-the-art denoising results.
@@ II Related Work
In this section, we offer a thorough review of HSI denoising, covering both model-based and deep learning-based methods. Additionally, we explore the application of transformers for HSI modeling.
@@ II-A Model-based HSI Denoising
Conventional model-based methods exploit sparse and low-rank representations to describe the physical properties of clean hyperspectral images (HSIs) for denoising [18, 12, 19, 20, 21, 22, 23]. For instance, BM4D [24], an extension of BM3D [24], utilizes collective sparse representation to capture similarity among nonlocal spatial-spectral cubes. Peng et al. [25] proposed tensor dictionary learning (TDL), which adaptively learns a dictionary from data to capture nonlocal spatial-spectral similarity. LLRT [26] incorporates an analysis-based hyper-Laplacian prior into the low-rank model, effectively restoring HSIs while preserving their spectral structure. NGMeet [27] uses low-rank Tucker decomposition and matrix rank minimization to capture global spectral correlation (GSC) and nonlocal self-similarity (NSS), respectively. FastHyDe [28] exploits spatial-spectral representations and their sparse, low-rank, and self-similarity characteristics in subspaces. Chen et al. [29] utilized total variation in the spectral subspace for fast HSI denoising. E-3DTV et al.[30] discerns global correlations and local differences across all bands by evaluating the sparsity of gradient maps in subspaces. Zhang [31] introduced a low-TR-rank regularizer and sparse regularizer to characterize the global low-rankness and sparsity of the transformed HSI, respectively, for denoising.
However, these model-based approaches typically involve iterative optimization and manual hyperparameter fine-tuning, making them computationally inefficient and possibly impractical for HSI denoising. Some methods have been proposed to combine model-based methods with data-driven learning [32, 33]. However, the inherent design of the model makes it difficult for these methods to meet the requirements in real-world scenarios where multiple types of noise are present.
@@ II-B CNN-based HSI Denoising
Since CNN showed its power on image denoising [34], deep learning has emerged as a critical approach in HSI denoising due to its robust learning capabilities [35]. Techniques utilizing 3-D convolution and its variations [36, 37, 11] have played a crucial role in capturing local spatial-spectral correlations for HSI denoising. Liu et al. [36] use a 3-D atrous CNN to expand the receptive field for HSI denoising, effectively handling mixed noise types while preserving image details. QRNN3D [11] integrates a recurrent unit after 3-D convolutions to leverage local spatial-spectral and global spectral correlation (GSC). NSSNN [38] combines an attention mechanism with the recurrent unit to extract GSC and non-local spatial features, achieving promising denoising performance. However, due to the limited long sequence modeling capability of CNNs, these methods often fail to capture the complete global spatial-spectral correlation in HSIs, leading to suboptimal exploitation of 3-D HSIs.
@@ II-C Transformer-based HSI Denoising
Transformers have shown remarkable success in various aspects of hyperspectral information processing, such as classification [39, 40, 41, 42, 43] and image recovering [44, 45, 46]. Hyperspectral imaging inherently involves 3-D data, where all dimensions exhibit strong correlations. While 3-D CNNs have traditionally been used for HSI modeling, the 3-D transformer [47] emerges as a promising alternative. However, suffering from the complexity of the self-attention mechanism, the 3-D transformer is often constrained since as the number of bands and image size increase, there is an explosion of space and computation. To address this issue, some methods use the transformer block to capture one of the long-range correlations. TRQ3D [13] incorporating a shifted windowing scheme to model nonlocal spatial relations and let RNNs model the spectral correlation. HSI-BERT [48] directly employs bidirectional encoder representations from transformers (BERTs) to encode semantic context-aware representations within each pixel, using multiple bands to capture spectral correlation. Furthermore, SST [9] utilizes a spatial transformer to model spatial correlation and a spectral transformer to model spectral correlation. However, limited by hardware conditions and the square growth of the transformer, we have always had difficulty in realizing global spatial-spectral correlation modeling in an all-in-one manner based on the transformer.
@@ II-D State Space Model
Recent advancements in the state space model (SSM) [15, 14] have made it widely applicable and opened up a new path for modeling sequentiality. Drawing on the theoretical basis of Classical SSM from control theory and leveraging the advantages of modern deep learning, emerging deep SSMs enable efficient computation in learning long-range dependencies over tens of thousands of timesteps. The field of computer vision has also seen an influx of SSM-based approaches [16, 49, 50, 51]. VmambaIR [52] introduced an SSM-based image restoration method. To fully exploit the 2-D image, VmambaIR incorporates a 2-D selective scan scheme to model the information flow in cross directions. However, the application of SSM to HSI denoising and the exploration of global spatial-spectral correlation in HSIs remain unexplored.
@@ III Method
In this section, we first introduce the preliminaries of SSM and Mamba. Then we provide a detailed exposition of the proposed SSUMamba, including the overall architecture, the Vision Mamba Block, the Spatial-Spectral Alternating Scan, and also the design details.
III-A Preliminaries: State Space Models
@@
As we can see Eq. (2) can be computed similarly to RNNs. To futher accelerate the computation, Gu et al [15] unroll the computation of the SSM into a convolution formulated as follows:
@@ III-B Overview Architecture
To model the global spatial-spectral correlation and enhance the 3-D characteristics of HSIs in denoising, we propose the SSUMamba, as shown in Fig. 1. Let be a clean HSI with pixels and bands. When subjected to additive noises , such as Gaussian noise, impulse noise, stripes, and deadlines, the observed HSI can be mathematically represented as:
@@
The Proposed SSUMamba is designed to learn the mapping from to for denoising in a U-shaped architecture, including encoder blocks, VMamba blocks with Spatial-Spectral Alternating Scan (SSAS), and decoder blocks. The encoder blocks employ downsampling and 3-D convolutions to extract multi-scale features from the observed HSI in multi-scale, while the VMamba blocks with SSAS are employed to model the global spatial-spectral correlation. The decoder blocks use upsampling and 3-D convolutions to reconstruct the clean HSI from the features extracted by the encoder blocks and the VMamba blocks with SSAS.
III-C Vision Mamba Block
@@
Mamba model [14] shows compelling characteristics in achieving global receptive fields, dynamic weights, and linear computational complexity. While the sequential nature of Mamba fits well with NLP tasks involving temporal data, it presents significant challenges when applied to vision data, which is non-sequential and rich in spatial information [50].
To address this issue, we introduce an adopted Mamba block for HSI modeling, as shown in Fig. 2. The Mamba block consists of Residual Blocks and a bidirectional Mamba layer. In our SSUMamba, the input HSIs is extracted by the encoder block into features with the shape of (B, H, W, D). Then, in the Mamba block, following [51], residual blocks with 3-D convolution, Instance Normalization, and Leaky ReLU are applied to the feature to model the global spatial-spectral correlation to enhance the local texture exploration. Moreover, the SSM is designed for the long-range correlation in unidirectional sequence data, while vision data is non-sequential. Thus, we applied a bidirectional Mamba layer [50] to model the information flow. In the bidirectional Mamba layer, the input feature is flattened into a sequence of vectors with the shape of (L, D), where , and then fed into the Mamba layer with two linear layers. In the first linear layer, the output feature is divided into two parts: one is processed in the forward direction, and the other is processed in the backward direction.
III-D Spatial-Spectral Alternating Scan
@@
The VMamba could save the modeling of SSM from unidirectional. However, in vision data, the correlation among pixels is in even more directions [50, 16]. For example, in 2-D vision data, 2D selective scan [52] is proposed to exploit the 2-D image in 4 directions: forward scan, backward scan, forward scan, and backward scan. Considering HSIs are 3-D data with strong spatial continuity and spectral correlation, we introduce a Spatial-Spectral Alternating Scan (SSAS) strategy, as show in Fig. 3, to fully exploit the information flow in all directions. The SSAS include 6 variants with different scanning priorities: , , , , , and . In each variant, the processing scans all pixels in the first direction and then proceeds to scan the next line in the second and third directions. We incorporate one of the variants of SSAS in the “Flatten” and “Reshape” operations in each Bidirectional Mamba to make the Mamba scan in a specific priority bidirectionally. Finally, in our SSUMamba, we use 6 VMamba blocks with all the 6 variants of SSAS to accomplish the modeling in all directions in HSIs.
III-E Architecture Design Details

### v3

@@ Abstract
Denoising hyperspectral images (HSIs) is a crucial preprocessing procedure due to the noise originating from intra-imaging mechanisms and environmental factors. Utilizing domain-specific knowledge of HSIs, such as spectral correlation, spatial self-similarity, and spatial-spectral correlation, is essential for deep learning-based denoising. Existing methods are often constrained by running time, space complexity, and computational complexity, employing strategies that explore these priors separately. While these strategies can avoid some redundant information, they inevitably overlook broader and more underlying long-range spatial-spectral information that positively impacts image restoration. This paper proposes a Spatial-Spectral Selective State Space Model-based U-shaped network, termed Spatial-Spectral U-Mamba (SSUMamba), for hyperspectral image denoising. We can obtain complete global spatial-spectral correlation within a module thanks to the linear space complexity in State Space Model (SSM) computations. We introduce a Spatial-Spectral Alternating Scan (SSAS) strategy for HSIs, which helps model the information flow in multiple directions in 3-D HSIs. Experimental results demonstrate that our method outperforms compared methods. The source code will be available at https://github.com/lronkitty/SSUMamba.
@@
Hyperspectral image denoising, selective state space model, mamba, global spatial-spectral correlation, deep learningI Introduction
@@
Deep learning-based denoising methods have recently garnered undivided attention. By leveraging data-driven learning, these methods can learn the 3-D nature of HSIs with spectral and spectral information. Due to the challenges posed by the design of methodology, running time, and spatial and computational complexity, existing deep learning-based methods often decompose this 3-D nature into several distinct domains for exploration, such as spatial self-similarity, spectral correlation, and spatial-spectral correlation. For example, recurrent neural networks (RNNs) and convolutional neural networks (CNNs) are resort to exploit spectral correlation and spatial-spectral correlation, respectively [11]. Moreover, attention mechanisms are concatenated to CNNs to capture spatial self-similarity further [12]. Transformer-based methods are also employed to capture long-range correlation in HSIs [9, 13, 13]. However, suffering from the complexity of the self-attention mechanism, these methods incorporate multiple modules to explore different domain knowledge, respectively, to reduce the length of tokens and avoid redundant information.
@@
However, given the tri-dimensional nature of HSIs, which demonstrate spatial solid continuity and spectral correlation, it is essential to explore the complete global spatial-spectral correlation, especially in image restoration tasks where detailed information tends to degrade. Broader and more underlying long-range spatial-spectral information has a positive impact that should be considered. The State Space Model (SSM) [14, 15], an approach rooted in control systems, has attracted attention for its linear complexity in handling long input sequences. Recently, Mamba [14], an optimization for SSM, has shown unique advantages in computer vision tasks [16, 17], while its linear complexity should offer promising advantages in global spatial-spectral correlation modeling in HSIs. Liu et al. [18] introduced a SSM-based method for HSI denoising to exploit the bidirectional spatial information in HSIs. However, how to apply the SSM to HSIs denoising to capture the global spatial-spectral correlation following the 3-D nature of HSIs remains an open question.
@@
This paper introduces a Spatial-Spectral Selective SSM-based HSI denoising method, Spatial-Spectral U-Mamba (SSUMamba), for HSI denoising. To address the disparity between image data and sequence data, we introduce the Spatial-Spectral Alternating Scan (SSAS) Mamba aimed to improve the exploration of local texture in the data and mitigate the unidirectional dependency issue present in the SSM. Our architecture incorporates a SSAS strategy for HSI data, which helps model the information flow in multiple directions in HSIs. Furthermore, we apply SSAS Mamba block in a U-shape network architecture to exploit multi-scale features and recover the HSI. Thanks to its linear complexity, SSUMamba can explore the complete global spatial-spectral correlation. The SSAS strategy further enhances the 3-D characteristics of HSI. Our SSUMamba demonstrates superior performance compared to other methods. The contributions of this article can be summarized as follows:
@@
To tackle the difference between HSI data and sequence data, we introduce the Spatial-Spectral Alternating Scan (SSAS) Mamba block to the HSI modeling to enhance local texture exploration and avoid unidirectional dependency in the SSM. Additionally, the SSAS strategy allows Mamba blocks to exploit information flows in all directions, enhancing the 3-D characteristics of HSIs.
Building upon the SSAS Mamba block, we propose a multi-scale SSM-based network, Spatial-Spectral U-Mamba (SSUMamba), for HSI denoising. Experiments demonstrate that SSUMamba achieves state-of-the-art denoising results.
@@ II Related Work
In this section, we offer a thorough review of HSI denoising, covering both model-based, CNN-based, and transformer-based methods. We also introduce the SSM in computer vision.
@@ II-A Model-based HSI Denoising
Conventional model-based methods exploit sparse and low-rank representations to describe the physical properties of clean hyperspectral images (HSIs) for denoising [19, 12, 20, 21, 22, 23, 24]. For example, BM4D [25], an extension of BM3D [25], utilizes collective sparse representation to capture similarity among nonlocal spatial-spectral cubes. Peng et al. [26] proposed tensor dictionary learning (TDL), which adaptively learns a dictionary from data to capture nonlocal spatial-spectral similarity. LLRT [27] incorporates an analysis-based hyper-Laplacian prior into the low-rank model, effectively restoring HSIs while preserving their spectral structure. NGMeet [28] uses low-rank Tucker decomposition and matrix rank minimization to capture global spectral correlation (GSC) and nonlocal self-similarity (NSS), respectively. FastHyDe [29] exploits spatial-spectral representations and their sparse, low-rank, and self-similarity characteristics in subspaces. Chen et al. [30] utilized total variation in the spectral subspace for fast HSI denoising. E-3DTV [31] discerns global correlations and local differences across all bands by evaluating the sparsity of gradient maps in subspaces. Zhang et al. [32] introduce a low-TR-rank regularizer and sparse regularizer to characterize the global low-rankness and sparsity of the transformed HSI, respectively, for denoising.
However, these model-based approaches typically involve iterative optimization and manual hyperparameter fine-tuning, making them computationally inefficient and possibly impractical for HSI denoising. Some methods have been proposed to combine model-based methods with data-driven learning [33, 34]. However, the inherent design for specifically noise types of the model makes it difficult for these methods to meet the requirements in real-world scenarios where multiple types of noise are present.
@@ II-B CNN-based HSI Denoising
Since CNN showed its power on image denoising [35], deep learning has emerged as a critical approach in HSI denoising due to its robust learning capabilities [36]. Techniques utilizing 3-D convolution and its variations [37, 38, 11] have played a crucial role in capturing local spatial-spectral correlations for HSI denoising. Liu et al. [37] use a 3-D atrous CNN to expand the receptive field for HSI denoising, effectively handling mixture noise types while preserving image details. QRNN3D [11] integrates a recurrent unit after 3-D convolutions to leverage local spatial-spectral and global spectral correlation (GSC). NSSNN [39] combines an attention mechanism with the recurrent unit to extract GSC and non-local spatial features, achieving promising denoising performance. However, due to the limited long sequence modeling capability of CNNs, these methods often fail to capture the complete global spatial-spectral correlation in HSIs, leading to suboptimal exploitation of 3-D HSIs.
@@ II-C Transformer-based HSI Denoising
Transformers have shown remarkable success in various aspects of hyperspectral information processing, such as classification [40, 41, 42, 43, 44] and image recovery [45, 46, 47]. Hyperspectral imaging inherently involves 3-D data, where all dimensions exhibit strong correlations. While 3-D CNNs have traditionally been used for HSI modeling, the 3-D transformer [48] emerges as a promising alternative. However, suffering from the complexity of the self-attention mechanism, the 3-D transformer is often constrained since as the number of bands and image size increase, there is an explosion of space and computation. To address this issue, some methods use the transformer block to capture one of the long-range correlations. TRQ3D [13] incorporates a shifted windowing scheme to model nonlocal spatial relations and let RNNs model the spectral correlation. HSI-BERT [49] directly employs bidirectional encoder representations from transformers (BERTs) to encode semantic context-aware representations within each pixel, using multiple bands to capture spectral correlation. Furthermore, SST [9] utilizes a spatial transformer to model spatial correlation and a spectral transformer to model spectral correlation. However, limited by hardware conditions and the square growth of the transformer, we have always had difficulty in realizing global spatial-spectral correlation modeling in an all-in-one manner based on the transformer.
@@ II-D State Space Model
Recent advancements in the state space model (SSM) [15, 14] have made it widely applicable and opened up a new path for modeling sequentiality. Drawing on the theoretical basis of Classical SSM from control theory and leveraging the advantages of modern deep learning, emerging deep SSMs enable efficient computation in learning long-range dependencies over tens of thousands of timesteps. The field of computer vision has also seen an influx of SSM-based approaches [16, 50, 51, 52]. Liu et al. [16] introduced a vision Mamba block with a 2-D selective scan scheme to model the information flow in cross directions. Shi et al. [53] further proposed an omni selective scan mechanism to model information flows to include the feature dimensions. SSM in HSI modeling has been widely explore [54, 55]. Liu et al. [18] apply the SSM to HSI denoising to exploit the bidirectional spatial information. However, the application of SSM to exploit the global spatial-spectral correlation in HSIs denoising remains unexplored.
@@ III Method
In this section, we first introduce the preliminaries of SSM and Mamba. Then, we provide a detailed exposition of the proposed SSUMamba, including the overall architecture, the Spatial-Spectral Alternating Scan (SSAS) Mamba block, and the design details.
III-A Preliminarie: State Space Models
@@
As we can see Eq. (2) can be computed similarly to RNNs. To further accelerate the computation, Gu et al [15] unroll the computation of the SSM into a convolution formulated as follows:
@@ III-B Overview Architecture
To model the global spatial-spectral correlation and enhance the 3-D characteristics of HSIs in denoising, we propose the SSUMamba, as shown in Fig. 1. Let be a clean HSI with pixels and B bands. When subjected to additive noises , such as Gaussian noise, impulse noise, stripes, and deadlines, the observed HSI can be mathematically represented as:
@@
The Proposed SSUMamba is designed to learn the mapping from to for denoising in a U-shaped architecture, including a feature extractor, encoder blocks, Spatial-Spectral Alternating Scan (SSAS) Mamba blocks, decoder blocks, a reconstructor. The feature extractor and reconstructor are utilized to extract HSI features with a shape of (B, H, W, D), where D represents the number of channels, and to recover the HSI from the features, respectively. The encoder blocks employ downsampling and 3-D convolutions to extract multi-scale features from the observed HSI feature in multi-scale, while the SSAS Mamba block are employed to model the local texture and global spatial-spectral correlation. The decoder blocks use upsampling and 3-D convolutions to reconstruct the features extracted by the encoder blocks and SSAS Mamba blocks.
III-C Spatial-Spectral Alternating Scan Mamba Block
@@
The Mamba model [14] shows compelling characteristics in achieving global receptive fields, dynamic weights, and linear computational complexity. While the sequential nature of Mamba fits well with NLP tasks involving temporal data, it presents significant challenges when applied to HSI data, which is non-sequential and has strong spatial continuity and spectral correlation. To address this issue, we introduce a Spatial-Spectral Alternating Scan (SSAS) Mamba block for HSI modeling, as shown in Fig. 2. The SSAS Mamba block consists of Residual Blocks and a SSAS Mamba layer.
Unlike the sequence data, vision data is rich in spatial information [51]. Additionally, HSIs feature strong local spatial-spectral correlation. In our SSAS Mamba Block, residual blocks with 3-D convolution, Instance Normalization, and Leaky ReLU are applied to the feature to model the local spatial-spectral correlation to enhance the local texture exploration in the SSAS Mama block.
@@
Moreover, the SSM models all tokens sequentially. However, in non-sequential vision data, the correlation among pixels is in more directions [51, 16]. For example, in vision data, a 2-D selective scan [16] is proposed to exploit the 2-D image in 4 directions: H forward scan, H backward scan, W forward scan, and W backward scan. Considering that HSIs are 3-D data with strong spatial continuity and spectral correlation, information flows are expected in both spatial and spectral dimensions to model global spatial-spectral correlation. Thus, we introduce a Spatial-Spectral Alternating Scan (SSAS) strategy, as shown in Fig. 3, to fully exploit the information flow in all directions. The SSAS include 6 variants with different scanning priorities: , , , , , and , where B, H, and W represent the band, height, and width of the HSI, respectively. The first 4 variants will pay attention to information in spatial dimension, while the last two variants focus on spectral correlation. In each variant, the processing scans all pixels in the first direction and then proceeds to scan the next line in the second and third directions. We incorporate one of the variants of SSAS after the residual blocks and flatten the input HSI feature with the shape of (B, H, W, D) into (L, D), where , to make the block scan in a specific priority. Finally, in our SSUMamba, we use 6 SSAS Mamba blocks with all the 6 variants of SSAS to accomplish the modeling in all directions in HSIs.
Thanks to SSAS, Mamba blocks can model the information flow in different priorities to model the 3-D characteristic of HSI. However, these information flows always sequence from the first pixel (1 in Fig. 3) to the last pixel (8 in Fig. 3). We further introduce a bidirectional Mamba layer to enhance the information flow. The feature from SSAS will be divided into two parts, one is processed in the forward direction, and the other is processed in the backward direction. After the bidirectional Mamba layer, the SSAS achieves the complete modeling of global spatial-spectral correlation.
III-D Architecture Design Details